OpenStudy (anonymous):

integrate intfrom 12, int from (1-x) to sqrt(x) x^2ydydx

OpenStudy (anonymous):

\[\int\limits_{1}^{1}\int\limits_{1-x}^{\sqrt{x}}x^2y dydx\]

OpenStudy (experimentx):

=0

OpenStudy (anonymous):

how do you figure

OpenStudy (experimentx):

1 to 1

OpenStudy (anonymous):

OOPS let me redo that lol

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\int\limits_{1-x}^{\sqrt{x}}x^2ydydx\]

OpenStudy (experimentx):

\[ \int_{1}^{2} y [x^3]_{1-x}^{\sqrt x} dy\]

OpenStudy (experimentx):

\[ [x^3/3]_{1-x}^{\sqrt x} * [y^2/2]_{1}^{2}\]